OpenStudy (anonymous):

Help please!! I'm lost on how to do this. A recursive function is shown below: F(1)=2 and f(n)=f(n-1)-5; n>1 Which of the following lists the terms in the sequence defined by this recursive function? A)2,3,8,13,18... B)2,7,12,17,22... C)2,-3,-8,-13,-18... D)2,-7,-12,-17,-22...

OpenStudy (anonymous):

clearly first term is 2 next term will be 2 -5 =-3 and third term will be -3-5=-8 and so on...

OpenStudy (anonymous):

Do you mind showing me how to do that with the formula? Thx so much for the help

jimthompson5910 (jim_thompson5910):

f(n)=f(n-1)-5 means "to find the next term, you take the last term and subtract 5" f(n) = nth term f(n-1) is the term just before the nth term

OpenStudy (anonymous):

Oh thanks! I didn't know about the last term.